OpenStudy (fambrinicrew):

5, 30, 155, 780, … 3905, 19530, 97655 905, 1020, 1145 1415, 2040, 2675 3065, 18690, 96815

OpenStudy (roxy.girl):

What is the Question?

OpenStudy (shootingstar28):

That is the question actually, I can help! :)

OpenStudy (fambrinicrew):

@Roxy.girl continue the pattern on the top with the mutiple choice below

OpenStudy (shootingstar28):

Okay so we are going to let x,y,z be and the next 3 numbers will be shown in the given sequence. We now construct a difference table as follows

OpenStudy (roxy.girl):

Do 780-155=?

OpenStudy (roxy.girl):

o.o What?

OpenStudy (roxy.girl):

We are doing two different things... xD

OpenStudy (shootingstar28):

Okay I'm gonna solve it simpler for ya :)

OpenStudy (shootingstar28):

So what we are going to do is multiply it by 5 the add 5

OpenStudy (shootingstar28):

Can you solve this? 5x5+5=30

OpenStudy (roxy.girl):

You need to just continue the pattern... You can do this simpler!

OpenStudy (shootingstar28):

30x5+5=155

OpenStudy (shootingstar28):

So here is what we have right now, 5x5+5=30 30x5+5=155

OpenStudy (shootingstar28):

Now 30x5+5=155 will become 155x5+5=780

OpenStudy (shootingstar28):

So as you conitue on you will get 780x5+5=3905 3905x5+5=19530 and Finally 19530x5+5=97655

OpenStudy (shootingstar28):

So: 5x5+5=30 30x5+5=155 155x5+5=780 780x5+5=3905 3905x5+5=19530 19530x5+5=97655

OpenStudy (shootingstar28):

The answer is 3905, 19530, 97655